HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

THE EXOMOL DATABASE: SPECTRLE

\underline{SERGEY N. YURCHENKO} , \underline{OLEKSIY SMOLA} , \underline{RYAN P. BRADY}, \underline{KYRIAKI KEFALA} , \underline{ANDREI SOKOLOV} , \underline{JINGXIN ZHANG} , \underline{GEORGI B. MITEV} , \underline{MARCO PEZZELLA} , \underline{ELIZABETH GUEST} ,

\underline{CHARLES BOWESMAN} , \underline{JONATHAN TENNYSON} , \underline{OLEG POLYANSKY} , Department of Physics and Astronomy, University College London, Gower Street,

WC1E 6BT London, United Kingdom.

ExoMol^{1} is a database of molecular line lists that can be used for spectral characterisation and simulation, and as input to atmospheric models of exoplanets, brown dwarfs and cool stars, and other models including those for combustion and sunspots.

ExoMol hosts over 80+ computed molecular line lists, where we present them as a game called ’spectrle’, a variation of the popular game ’wordle’, where one has to guess what molecule belongs to each presented spectrum. A prize will be given to the person who guesses the highest number of molecular spectra.

We present an update to the ExoMol database, featuring new line lists for AlCl^{2}, LiOH^{3}, SO^{4}, H_{2}CS^{5}, H_{3}^{+}^{6}, CaOH^{7}, SiN^{8}, NaO^{9} and updated line lists for CaH, MgH^{10}, AlO^{11}, NO^{12}.

1https://www.exomol.com/

2https://doi.org/10.1093/mnras/stac3757 SN Yurchenko, E Nogué, AAA Azzam, J Tennyson - Monthly Notices of the Royal Astronomical Society, 2023

3https://www.exomol.com/data/molecules/LiOH/7Li-16O-1H/OYT7 Owens A., Pavlenko, Ya., Mitrushchenkov, A., Koput, J. Yurchenko, S.N., Tennyson J., ”ExoMol line lists - LI. Molecular line list for lithium hydroxide (LiOH)”, Monthly Notices of the Royal Astronomical Society Submitted (2023).

4doi:10.1039/D2CP03051AR. P. Brady, S. N. Yurchenko, G.-S. Kim, W. Somogyi and J. Tennyson, An ab initio study of the rovibronic spectrum of sulphur monoxide (SO): diabatic vs. adiabatic representation, Phys. Chem. Chem. Phys., 24, 24076 (2022)

5https://doi.org/10.1093/mnras/stad111 T Mellor, A Owens, J Tennyson, SN Yurchenko - Monthly Notices of the Royal Astronomical Society, 2023

6https://doi.org/10.1093/mnras/stad050 CA Bowesman, II Mizus, NF Zobov, OL Polyansky - Monthly Notices of the Royal Astronomical Society, 2023

7https://doi.org/10.1093/mnras/stac2462 A Owens, A Mitrushchenkov, SN Yurchenko - Monthly Notices of the Royal Astronomical Society, 2022

8https://doi.org/10.1093/mnras/stac2004 M Semenov, N Clark, SN Yurchenko, GS Kim - Monthly Notices of the Royal Astronomical Society, 2022

9https://doi.org/10.1093/mnras/stab3357 GB Mitev, S Taylor, J Tennyson, SN Yurchenko - Monthly Notices of the Royal Astronomical Society, 2022

10https://doi.org/10.1093/mnras/stac371 A Owens, S Dooley, L McLaughlin, B Tan, G Zhang - Monthly Notices of the Royal Astronomical Society, 2022

11https://doi.org/10.1093/mnras/stab2525 CA Bowesman, M Shuai, SN Yurchenko, J Tennyson - Monthly Notices of the Royal Astronomical Society, 2021

12https://doi.org/10.1093/mnras/stab1154 Q Qu, SN Yurchenko, J Tennyson - Monthly Notices of the Royal Astronomical Society, 2021

p-number: p213  Submitted on Fri, 16 Jun 2023 00:43:59 +0200